STAR Protected

TPC weekly

2021-08-18 15:00
America/New York
Wednesday, 18 August 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-08-04 15:00
America/New York
Wednesday, 4 August 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-08-11 15:00
America/New York
Wednesday, 11 August 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-07-21 15:00
America/New York
Wednesday, 21 July 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-07-28 15:00
America/New York
Wednesday, 28 July 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-07-14 15:00
America/New York
Wednesday, 14 July 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-07-07 15:00
America/New York
Wednesday, 7 July 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-06-30 15:00
America/New York
Wednesday, 30 June 2021
virtual , at 19:00 (GMT), duration : 00:00
TPC meeting on Wednesday 3.00pm

agenda
- hardware status
- operation
   failures, humidity, magnet current measurement by C-AD
   additional FF cosmic?
- software
 - calibration status
- plans for shutdown (reminder)





Topic: TPC weekly
Time: Jun 23, 2021 03:00 PM Eastern Time (US and Canada)
        Every week on Wed, until Aug 4, 2021, 7 occurrence(s)

        Jun 30, 2021 03:00 PM
        Jul 7, 2021 03:00 PM
        Jul 14, 2021 03:00 PM
        Jul 21, 2021 03:00 PM
        Jul 28, 2021 03:00 PM
        Aug 4, 2021 03:00 PM
Please download and import the following iCalendar (.ics) files to your calendar system.
Weekly: https://bnl.zoomgov.com/meeting/vJItdOmurDksH_3czqVgkBAd421P_ZMYLYo/ics?icsToken=98tyKuiurjgqHNaXthzBeI89EqPga-nFiWJmlfBehg7TBCtxQweiZLNyDYhSKfTp

Join ZoomGov Meeting
https://bnl.zoomgov.com/j/1603062265?pwd=Tm9Nb3lpcWt0L1ZUcUg0TjV0T2Vydz09

Meeting ID: 160 306 2265
Passcode: 880354
One tap mobile
+16692545252,,1603062265#,,,,*880354# US (San Jose)
+16468287666,,1603062265#,,,,*880354# US (New York)

Dial by your location
        +1 669 254 5252 US (San Jose)
        +1 646 828 7666 US (New York)
        +1 669 216 1590 US (San Jose)
        +1 551 285 1373 US
Meeting ID: 160 306 2265
Passcode: 880354
Find your local number: https://bnl.zoomgov.com/u/aM2ZS218b

Join by SIP
1603062265@sip.zoomgov.com

Join by H.323
161.199.138.10 (US West)
161.199.136.10 (US East)
Meeting ID: 160 306 2265
Passcode: 880354

TPC weekly

2021-06-23 15:00
America/New York
Wednesday, 23 June 2021
virtual , at 19:00 (GMT), duration : 00:00
TPC meeting on Wedbesday 3.00pm

agenda
- hardware status
- operation
   failures, humidity, magnet current measure
- software
 - calibration status
- plans for shutdown (reminder)





Topic: TPC weekly
Time: Jun 23, 2021 03:00 PM Eastern Time (US and Canada)
        Every week on Wed, until Aug 4, 2021, 7 occurrence(s)
        Jun 23, 2021 03:00 PM
        Jun 30, 2021 03:00 PM
        Jul 7, 2021 03:00 PM
        Jul 14, 2021 03:00 PM
        Jul 21, 2021 03:00 PM
        Jul 28, 2021 03:00 PM
        Aug 4, 2021 03:00 PM
Please download and import the following iCalendar (.ics) files to your calendar system.
Weekly: https://bnl.zoomgov.com/meeting/vJItdOmurDksH_3czqVgkBAd421P_ZMYLYo/ics?icsToken=98tyKuiurjgqHNaXthzBeI89EqPga-nFiWJmlfBehg7TBCtxQweiZLNyDYhSKfTp

Join ZoomGov Meeting
https://bnl.zoomgov.com/j/1603062265?pwd=Tm9Nb3lpcWt0L1ZUcUg0TjV0T2Vydz09

Meeting ID: 160 306 2265
Passcode: 880354
One tap mobile
+16692545252,,1603062265#,,,,*880354# US (San Jose)
+16468287666,,1603062265#,,,,*880354# US (New York)

Dial by your location
        +1 669 254 5252 US (San Jose)
        +1 646 828 7666 US (New York)
        +1 669 216 1590 US (San Jose)
        +1 551 285 1373 US
Meeting ID: 160 306 2265
Passcode: 880354
Find your local number: https://bnl.zoomgov.com/u/aM2ZS218b

Join by SIP
1603062265@sip.zoomgov.com

Join by H.323
161.199.138.10 (US West)
161.199.136.10 (US East)
Meeting ID: 160 306 2265
Passcode: 880354


TimeTalkPresenter
15:00dE/dx from express production 3p85GeV_fixedTarget 2021 ( 00:05 ) 0 filesYuri Fisyak

TPC weekly

2021-06-16 15:00
America/New York
Wednesday, 16 June 2021
virtual , at 19:00 (GMT), duration : 00:00
Minutes TPC meeting Wed 15

Present:
Prashanth, Gene, Alexei, Babu, Dan, Ben,Yuri, Hongwei,Flemming
Hardware:
Two major issues.
-    West laser is unstable, and difficult to tune. Probably not mirror issues.
-    Sector 15 HV section 3 (itpc) tripped yesterday afternoon (June 15) and cannot hold voltage. Turned to 0 overnight. Today it held 1035 for 10 min but then tripped. AL left it at 500 V.
-    Alexei will investigate further on Monday’s access day.

•    Since last week 21-4 is now masked out. Tonko report in his records
•    iS21-4 dead [11 Jun]
o    comes up but then looks like it loses power and dies again
o    perhaps a FEE is shorting out?
o    Masked
S24-4 masked out for reminder of run.
Included the bad HV section about 6-7% of iTPC is out

Run-20 calibration (Babu). Has problem with failed jobs. Not enough jobs has completed to finish period 1 dynamic. Will discuss failures with Gene.

Yuri reported no progress on FF calibrations and understanding.

Reports that dE/dx for the FXt is down ~8% compared to 7.7 GeV where it was first calibrated, and additional drift down by about 4% here in June.
Gene reports that for run 22 there are ideas to increase rates, reduce data volume by masking out ~10 innermost row of iTPC. These are ‘black in pp 500 anyhow. Gene will run some simple simulations (Jim Thomas macro’s) to see what effect in tracking will be w/o these. Full simulation may also be needed. There will be a talk at the analysis meeting on this by Tonko.

Hongwei confirmed that the discrepancy between HLT good and offline is about 3-4% after dealing properly with vertex selection in his offline runs.

TPC weekly

2021-06-09 15:00
America/New York
Wednesday, 9 June 2021
virtual , at 19:00 (GMT), duration : 00:00
 agenda:
- operation
- calibration status
- FF field

Yuri posted additional slide on FF DCA_y
https://drupal.star.bnl.gov/STAR/system/files/OO200GeV2021RFvsFF_0.pdf


Tpc meeting June 9, 2021
Present: Gene, Yuri, Dan, Babu, Alexei, HongWei, Flemming

Hardware:
Humidity has been creeping up a bit. Yesterday some high current in fieldcage due to humidity.
Quote from Gene
“I find that the blower A/C being off on June 5th did raise the inner field cage currents during two physics runs:
22156024 and 026

I propose simply to mark these two runs as bad. I would advise in situations like this to stop the fill and address the blower ASAP rather than waiting for RHIC operators to reach their own end-of-fill.


The number of auto-recovery has gone up during the 3.985 GeV FXT . Beam is likely dirtier. Date for 24:s4 looks fine

Operation:
The ops crew noted alarm for magnet current at ops meeting. It seems to be correlated to the lower mean current observed the last day. This was observed earlier in the run by Yuri  see 
https://drupal.star.bnl.gov/STAR/blog/fisyak/averaged-star-magnet-current-table. The large oscillation is due to measurements  (+-2A)– this is understood. The drop today as well as drops earlier in the run is not. The oscillation gives a K0S mass change of ~0.4 GeV/c**2.
Yuri was asked to look at mean of K0S mass for the periods (or a period) where the mean is down to -4112 for average to identify if it is due to measurement or to change in magnetic field.
Alexei has earlier noted that for next year we should have a NMR installed with high precision field measurement.

Calibration:
At this point the TPC calibration are done for run-19.
Babu plans to have an update next week for the run-20 iterations.
Yuri had an additional plot of the FF local DCA_y  ivs. Sector number
https://drupal.star.bnl.gov/STAR/system/files/OO200GeV2021RFvsFF_0.pdf
The fits correspond to a rotation around the vertical axis of the field relative to TPC, and is of similar magnitude as the measured difference between field and TPC at zero-filed and incorporated in the ‘twist distortion correction. Yuri has discussed with Alexei and Bill C, to see if this can be  understood i.e why changing the field to result in such change.

Photon v2 Analysis

iTPC sector 24

sector 24.4 issues during day 143-144
Unfortunately managed to delete conent when trying to insert table of runs with this issue.
Will restore later.  (5/28)

TPC weekly

2021-05-12 15:00
America/New York
Wednesday, 12 May 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC meeting May 12, 2021

 

Present: Alexei, Gene, Yuri, Hongwei and Flemming

 

Hardware:

Basically in good shape. Struggled with HV GUI’s likely due to ITD scanning during the access day. Sector 20 outer was repaired by Tonko and Mike, bad RORC card.

Has observed the part short of west filed cage.

 

Alexei has the calibrated NMR in house.

 

Operations:

The O+O running  very clean. The 1.5mrad crossing angle and levelling at ~3-4kHz Mbias results in very little if any dynamic space charge.

This  bodes well for the FF setting to be run in about a week.

For the central collisions to be run right after the RFF 400 M Mbias the Mbias rate will be possibly ~20kHz.

 

Calibrations:

 

Gene looked at h+/hi ratios vs. pt for the 19.6 Gev au+au and reports seeing sector to sector variations;

https://drupal.star.bnl.gov/STAR/blog/genevb/h-h-run-19-196-gev

These variations are at similar level as after run 12 alignment calibrations i.e little overall effect but sector to sector variations.

The sectors showing largest effect correlates with Zaochen's findings in mass2 splitting in BTOF

Thus the variation is most likely due to TPC calibrations.

Yuri commented that until we have the FF data  its unlikely that underlying reason, and improved calibration can be made. It is also likely to be a large effort, not a few weeks.

 

Yuri reported that dE/dx calibration for 14.5 and 19.6 is completed

It was also reported to S&C management

https://drupal.star.bnl.gov/STAR/system/files/19%20and%2014p5%20GeV%20dEdx%20Calibration.pdf

TPC weekly

2021-05-19 15:00
America/New York
Wednesday, 19 May 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-05-26 15:00
America/New York
Wednesday, 26 May 2021
virtual , at 19:00 (GMT), duration : 00:00
agenda
- hardware status
    Gene: humidity impacts seen on inner field cage currents. See image shown further down, below agenda. More on this topic in general in my blog.
- operations
- calibration status

Flemming Videbaek is inviting you to a scheduled ZoomGov meeting.
 
Topic: TPC weekly
 
        May 26, 2021 03:00 PM
        Jun 2, 2021 03:00 PM
        Jun 9, 2021 03:00 PM
Please download and import the following iCalendar (.ics) files to your calendar system.
Weekly: https://bnl.zoomgov.com/meeting/vJIsd-utqz0oGdOZMvWhDIufxGjndzII97w/ics?icsToken=98tyKuiurzsoH9GTshrBeI86EsH4aur70n5proxnsQnSCwR6dwXuEehMNqhWOYGH
 
Join ZoomGov Meeting
https://bnl.zoomgov.com/j/1610255623?pwd=N2FmemwvM3prQzExYnc0VlptdVpGUT09
 
Meeting ID: 161 025 5623
Passcode: 629521
One tap mobile
+16692545252,,1610255623#,,,,*629521# US (San Jose)
+16468287666,,1610255623#,,,,*629521# US (New York)
 
Dial by your location
        +1 669 254 5252 US (San Jose)
        +1 646 828 7666 US (New York)
        +1 669 216 1590 US (San Jose)
        +1 551 285 1373 US
Meeting ID: 161 025 5623
Passcode: 629521
Find your local number: https://bnl.zoomgov.com/u/ac5iOCm2Y1
 
Join by SIP
1610255623@sip.zoomgov.com
 
Join by H.323
161.199.138.10 (US West)
161.199.136.10 (US East)
Meeting ID: 161 025 5623
Passcode: 629521
 
TimeTalkPresenter
14:00OO 200 GeV 2021 RF vs FF ( 00:20 ) 1 file

TPC weekly

2021-06-02 15:00
America/New York
Wednesday, 2 June 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC weekly

2021-04-14 15:00
America/New York
Wednesday, 14 April 2021
virtual , at 19:00 (GMT), duration : 00:00
Agenda
- hardware update
- operations
- calibration updates


       Apr 14, 2021 03:00 PM
        Apr 21, 2021 03:00 PM
Please download and import the following iCalendar (.ics) files to your calendar system.
Weekly: https://bnl.zoomgov.com/meeting/vJIsf--vqzMqHOQ-n7SvocdwW3KiZSeaERo/ics?icsToken=98tyKuiurzMsHdGdsB_BeI86EMGgaOr7r2BmhZIQkkjWFS9QZhTBZclLCIF6Ef3i
 
Join ZoomGov Meeting
https://bnl.zoomgov.com/j/1618675806?pwd=d2Njdjg4WHhMRzU4OEVmY3dSRktKUT09
 
Meeting ID: 161 867 5806
Passcode: 514827
One tap mobile
+16692545252,,1618675806#,,,,*514827# US (San Jose)
+16468287666,,1618675806#,,,,*514827# US (New York)
 
Dial by your location
        +1 669 254 5252 US (San Jose)
        +1 646 828 7666 US (New York)
        +1 551 285 1373 US
        +1 669 216 1590 US (San Jose)
Meeting ID: 161 867 5806
Passcode: 514827
 

TPC weekly

2021-04-21 15:00
America/New York
Wednesday, 21 April 2021
virtual , at 19:00 (GMT), duration : 00:00

TPC meeting April21 minutes

 

Present:

Alexei,Babu, Flemming,Gene,Hongwei,Yuri

 

TPC running well. The pizo driver for the laser failed , no power,was adjusted during quick ccess, and then rebooted today during access. Lost about one laser run during. Will investigate how remote reboot can be done. Such failure has happened 1-2 per run.

 

Flemming present a few more observation on the single bad pad hot due to values this seem to have pedestal subtracted incorrectly. This is intermittent and usually, but not always clears on next run. Also it only 5-10% of events with this failure. See  https://drupal.star.bnl.gov/STAR/blog/videbaks/tpc-qa-2021
for more details

 

The 19.6 GeV calibration sample is complete. Includes dynamic space charge, and event by event t0 correction. EPD had time offsets calibrated/changed at some point in time during the 19.6 GeV data sets. This change is now in DB.  The 14.5 GeV  should be done by the weekend. Babu will start looking at the run 20 soon.

 

The sample(s) can now be checked for BTOF calibrations, and dE/dx.

If the mass square difference seen in east persists with the DbTpcMaker that was updated, there will be additional work to identify this issue. The TOF group will not sign off with a mass split of approx. 1%.

 

 

TPC weekly

2021-04-28 15:00
America/New York
Wednesday, 28 April 2021
virtual , at 19:00 (GMT), duration : 00:00
agenda
- hardware status
- operations
- calibration status

Flemming Videbaek is inviting you to a scheduled ZoomGov meeting.
 
Topic: TPC weekly
Time: Apr 28, 2021 03:00 PM Eastern Time (US and Canada)
        Every week on Wed, until Jun 9, 2021, 7 occurrence(s)
        Apr 28, 2021 03:00 PM
        May 5, 2021 03:00 PM
        May 12, 2021 03:00 PM
        May 19, 2021 03:00 PM
        May 26, 2021 03:00 PM
        Jun 2, 2021 03:00 PM
        Jun 9, 2021 03:00 PM
Please download and import the following iCalendar (.ics) files to your calendar system.
Weekly: https://bnl.zoomgov.com/meeting/vJIsd-utqz0oGdOZMvWhDIufxGjndzII97w/ics?icsToken=98tyKuiurzsoH9GTshrBeI86EsH4aur70n5proxnsQnSCwR6dwXuEehMNqhWOYGH
 
Join ZoomGov Meeting
https://bnl.zoomgov.com/j/1610255623?pwd=N2FmemwvM3prQzExYnc0VlptdVpGUT09
 
Meeting ID: 161 025 5623
Passcode: 629521
One tap mobile
+16692545252,,1610255623#,,,,*629521# US (San Jose)
+16468287666,,1610255623#,,,,*629521# US (New York)
 
Dial by your location
        +1 669 254 5252 US (San Jose)
        +1 646 828 7666 US (New York)
        +1 669 216 1590 US (San Jose)
        +1 551 285 1373 US
Meeting ID: 161 025 5623
Passcode: 629521
Find your local number: https://bnl.zoomgov.com/u/ac5iOCm2Y1
 
Join by SIP
1610255623@sip.zoomgov.com
 
Join by H.323
161.199.138.10 (US West)
161.199.136.10 (US East)
Meeting ID: 161 025 5623
Passcode: 629521
 
 
TimeTalkPresenter
15:00Averaged STAR magnet current table ( 00:10 ) 0 files
15:00Averaged STAR magnet current table ( 00:10 ) 0 files
15:102nd round with Run XIX calibration production for 19GeV and 14p5GeV samples ( 00:20 ) 0 filesYuri Fisyak